# Maximal discs of Weil-Petersson class in $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$

### v1

@@ 2.2. Weil-Petersson universal Teichmüller space
where $d^{2}z=dxdy$ for $z=x+\sqrt{-1}y$. Denote by $\mathcal{O}(\mathbb{D})_{1}$ the subgroup of $L^{\infty}(\mathbb{D})_{1}$ generated by $\mu\in\Omega^{-1,1}(\mathbb{D})$ with $||\mu||_{\infty}<\delta$ where $\delta$ is a postive real number satisfying condition in Corollary 2.6 in [15]. For every $\mu\in\mathcal{O}(\mathbb{D})_{1}$, let $V_{\mu}\subset U_{\mu}\subset T(1)$ be the image under the map $h_{\mu}^{-1}=\mathrm{P}\circ R_{\mu}\circ\Lambda$ of the open ball of radius $\sqrt{\pi/3}$ about the origin in $A_{2}(\mathbb{D}^{*})$. Let $\tilde{h}_{\mu}=h_{\mu}\Big{|}_{V_{\mu}}:V_{\mu}\to A_{2}(\mathbb{D^{*}})$. Now we consider the covering
@@
By Theorem 2.10 in [15], the above covering endows $T(1)$ with the structure of a complex Hilbert manifold modeled on the Hilbert space $A_{2}(\mathbb{D}^{*})$. But, the space $T(1)$ is not connected with respect to the topology induced by the Hilbert manifold structure.
@@
is a biholomorphic mapping of Hilbert manifolds. Hence the distribution $\mathcal{D}_{T}$ on $T(1)$ is integrable. Their integral manifolds correspond to the components $(\phi+A_{2}(\mathbb{D}^{*}))\cap\beta(T(1))$. For every $[\mu]\in T(1)$, we denote by $T_{[\mu]}(1)$ the component of the Hilbert manifold $T(1)$ containing $[\mu]$. The Hilbert manifold $T_{[\mu]}(1)$ is the integral manifold of the distribution $\mathcal{D}_{T}$ passing through $[\mu]\in T(1)$. In particular, the component of the origin $0\in T(1)$ is denoted by $T_{0}(1)$, and is called Weil-Petersson universal Teichmüller space.
The Weil-Petersson metric on the Hilbert manifold $T_{0}(1)$ is a Hermitian metric defined by the Hilbert space inner product on tangent space, which is identified with the Hilbert space $H^{-1,1}(\mathbb{D})$ by a right translation. Thus the Weil-Petersson metric is a right invariant metric on $T_{0}(1)$ defined at the origin of $T_{0}(1)$ by
@@
To every $\mu\in H^{-1,1}(\mathbb{D})$, there corresponds a vector field $\frac{\partial}{\partial\epsilon_{\mu}}$ over $V_{0}$. We set for every $\kappa\in V_{0}$,
@@ 3. Maximal discs of Weil-Petersson class
In this section, we introduce and study maximal discs of Weil-Petersson class in the 3-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. The deformation space of maximal discs of Weil-Petersson class is understood as the cotangent bundle $T^{*}T_{0}(1)$ of the Weil-Petersson universal Teichmüller space $T_{0}(1)$, and we show that this can be identified with $T_{0}(1)\times T_{0}(1)$ by the map originally defined by Mess in [10]. We refer to the appendix of this paper for the basic terminology related to $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ and to [5] for the extensive introduction for the Anti de Sitter geometry.
@@ 3.1. Maximal discs and Gauss maps
For an Beltrami differential $\mu$ representing an element in the universal Teichmüller space $T(1)$, let us denote by $(\mathbb{D},z)$ the corresponding unit disc with the coordinate map $z$. Note that $z$ is given by the quasi-conformal map $w_{\mu}:\mathbb{D}\to\mathbb{D}$ introduced in the subsection 2.1. Let us also fix the conformal structure on the standard disc $\mathbb{D}\subset\mathbb{C}$ and denote it by $(\mathbb{D},w)$ with its coordinate map $w$.
For a given $(\mathbb{D},z)$, let us denote the spacelike conformal embedding of $(\mathbb{D},z)$ into the 3-dimensional anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ by
$\sigma:(\mathbb{D},z)\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}.$
For a spacelike conformal embedding $\sigma$, the image $\Sigma:=\sigma(\mathbb{D})$ is called maximal if the mean curvature $H_{\sigma}$ identically vanishes on $\Sigma$.
By Proposition 3.1 and Theorem 3.3 of [2], For the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, each component of the Gauss map $G_{\pm}:\Sigma\to(\mathbb{D},w)$ is a harmonic map, that is, satisfies the following equality
(3.1) $(G_{\pm})_{z\bar{z}}+(\psi_{w}\circ G_{\pm})(G_{\pm})_{z}(G_{\pm})_{\bar{z}}=0,$
where $e^{\psi(w)}|dw|^{2}$ denotes the hyperbolic metric on $(\mathbb{D},w)$. The first fundamental form $I$, that is, the induced metric on $\Sigma$ is given by
(3.2) $I=e^{\phi}|dz|^{2}:=e^{\psi\circ G_{\pm}}\big{|}(G_{\pm})_{z}\big{|}^{2}|dz|^{2}.$
The Hopf differential $\Phi(G_{\pm})$ of the harmonic map $G_{\pm}:\Sigma\to(\mathbb{D},w)$ satisfies
(3.3) $\Phi(G_{+})=e^{\psi\circ G_{+}}(G_{+})_{z}(\overline{G_{+}})_{z}\,dz^{2}=-\Phi(G_{-})=-e^{\psi\circ G_{-}}(G_{-})_{z}(\overline{G_{-}})_{z}\,dz^{2},$
and the second fundamental forms of $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is given by
@@
We refer to the section 3 of [2] and the section 5 of [7] for more details about these equalities.
By the Gauss equation for $\Sigma$, we have
@@
Note that (3.5) can also be derived from the harmonic map equation for $G_{\pm}$ given in (3.1). We also have
@@
The Gaussian curvature $K_{\phi}$ for the metric $e^{\phi}|dz|^{2}$ on $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is given by
@@
From (3.2) and (3.1), denoting by $G=G_{\pm}$,
@@
$\phi_{z\bar{z}}=(\psi_{w\bar{w}}\circ G)\big{(}|G_{z}|^{2}-|G_{\bar{z}}|^{2}\big{)}=\frac{1}{2}e^{\psi\circ G}\big{(}|G_{z}|^{2}-|G_{\bar{z}}|^{2}\big{)}.$
Here the second equality follows from the Liouville equation for $e^{\psi}$, that is, $\psi_{w\bar{w}}=1/2e^{\psi}$. Hence, we have
@@
By the Gauss equation for $\Sigma$, $K_{\phi}=-1+\kappa^{2}$ where ±$\pm\kappa$ denote the principal curvatures of the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Hence,
@@
For the hyperbolic metric on $(\mathbb{D},w)$, its pullback by $G_{\pm}$ determines a new conformal structure on $\Sigma$ respectively. We denote them by $(\mathbb{D},z_{\pm})$ respectively. By Theorem 3.2 of [18], for the Hopf differential ±$\pm\Phi=\Phi(G_{\pm})$ on $\Sigma\cong(\mathbb{D},z)$ and $(\mathbb{D},z_{\pm})$, there is a harmonic map $F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm})$ such that the associated Hopf differential is ±$\pm\Phi$. Moreover, we have
@@
(3.8) $e^{\phi}=e^{\psi\circ G_{\pm}}|(G_{\pm})_{z}|^{2}=e^{\psi\circ F_{\pm}}|(F_{\pm})_{z}|^{2},\qquad\mu_{G_{\pm}}=\mu_{F_{\pm}},$
where $e^{\psi}$ denotes the hyperbolic metric density on $(\mathbb{D},w)$, $(\mathbb{D},z_{\pm})$ respectively.
Proof.
Recall that the harmonic map condition of $G_{\pm}$, $F_{\pm}$ depends on the conformal structure on $\Sigma$ rather than the metic on $\Sigma$, and that the concerned quantities of $G_{\pm}$, $F_{\pm}$ are the square of the holomorphic energy densities times the hyperbolic metric density on $\Sigma$, that is,
$e^{\psi\circ G_{\pm}}|(G_{\pm})_{z}|^{2}=|\partial G_{\pm}|^{2}e^{\psi(z)},\qquad e^{\psi\circ F_{\pm}}|(F_{\pm})_{z}|^{2}=|\partial F_{\pm}|^{2}e^{\psi(z)}.$
Then, by Theorem 2.1 of [18] (see also the section 4 of [22]), the functions
(3.9) $\mathfrak{w}(G_{\pm}):=1/2\log|\partial G_{\pm}|^{2},\qquad\mathfrak{w}(F_{\pm}):=1/2\log|\partial F_{\pm}|^{2}$
satisfy the following equation
$14e^{-\psi}\mathfrak{w}_{z\bar{z}}=-(K\circ u)\big{(}e^{2\mathfrak{w}}-e^{-2\mathfrak{w}}e^{-2\psi}|\Phi|^{2}\big{)}-$ 1
where $K$ denotes the Gaussian curvature of the target disc of $G_{\pm}$ or $F_{\pm}$ and the map $u$ is $G_{\pm}$ or $F_{\pm}$ depending on the case. But, $K(u)\equiv-1$ for these cases. Then, by Corollary 8 of [22], the solution of this equation with $K\circ u\equiv-1$ is unique. Hence, $\mathfrak{w}(G_{+})=\mathfrak{w}(G_{-})=\mathfrak{w}(F_{+})=\mathfrak{w}(F_{-})$, which implies the first equality of $e^{\psi\circ G_{\pm}}|(G_{\pm})_{z}|^{2}=e^{\psi\circ F_{\pm}}|(F_{\pm})_{z}|^{2}$ in (3.8). Note that the Beltrami differentials $\mu_{G_{\pm}}$ and $\mu_{F_{\pm}}$ satisfy
$\mu_{G_{\pm}}=\pm\overline{\Phi}e^{-\phi}=\mu_{F_{\pm}},$
whhich implies the second equality in (3.8). ∎
By Proposition 3.2, $(F_{+},F_{-})$ shares some basic properties with the Gauss map $(G_{+},G_{-})$, and we call $(F_{+},F_{-})$ the induced Gauss map.
Combining the above constructions for two cases of $G_{\pm}$, we have the following commutative diagram
@@
Here $z,z_{+},z_{-}$ denote the quasi-conformal maps from $(\mathbb{D},w)$ to $(\mathbb{D},z)$, $(\mathbb{D},z_{+})$, $(\mathbb{D},z_{-})$ respectively. We also abuse notations by using the same notations $F_{+}$, $F_{-}$ for the central horizontal maps, which are obviously induced by $F_{\pm}:\Sigma\to(\mathbb{D},z_{\pm})$. We may also assume that all the maps $z,z_{\pm}$ and $F_{\pm}$ are normalized, that is, preserves $1,-1,-\sqrt{-1}$.
Remark 3.3.
By the construction of the harmonic map $F_{\pm}:(\mathbb{D},z)\to(\mathbb{D},z_{\pm})$, we have $F_{\pm}\circ z=z_{\pm}$ as explained above. Hence, the harmonic map $F_{\pm}$ is compatible with the coordinate maps $z,z_{\pm}$ in the above sense. But, such a property does not holds for $G_{\pm}:(\mathbb{D},z)\to(\mathbb{D},w)$, that is, $G_{\pm}\circ z\neq\mathrm{id}$.
By Theorem 4.1 of [2], Theorem 4.1 of [4], and the previous construction, we have
@@
There are one-to-one correspondences between the followings:
@@
    a maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ such that the Gaussian curvature of the induced metric is negative,
@@
    an orientation preserving minimal diffeomorphism $G:(\mathbb{D},w)\to(\mathbb{D},w)$ where $(\mathbb{D},w)$ is equipped with the hyperbolic metric,
@@
    an orientation preserving minimal diffeomorphism $F:(\mathbb{D},z_{+})\to(\mathbb{D},z_{-})$ where $(\mathbb{D},z_{\pm})$ is equipped with the hyperbolic metric,

### v2

@@ 2.2. Weil-Petersson universal Teichmüller space
where $d^{2}z=dxdy$ for $z=x+\sqrt{-1}y$. Let $\mathcal{O}(\mathbb{D})_{1}$ denote the subgroup of $L^{\infty}(\mathbb{D})_{1}$ generated by $\mu\in\Omega^{-1,1}(\mathbb{D})$ with $||\mu||_{\infty}<\delta$ where $\delta$ is a positive real number satisfying condition in Corollary 2.6 in [18]. For each $\mu\in\mathcal{O}(\mathbb{D})_{1}$, let $V_{\mu}\subset U_{\mu}\subset T(1)$ be the image under the map $h_{\mu}^{-1}=\mathrm{P}\circ R_{\mu}\circ\Lambda$ of the open ball of radius $\sqrt{\pi/3}$ centered at the origin in $A_{2}(\mathbb{D}^{*})$. Define
(2.14) $\tilde{h}_{\mu}=h_{\mu}\Big{|}_{V_{\mu}}:V_{\mu}\to A_{2}(\mathbb{D^{*}}).$
Now, consider the covering
@@
By Theorem 2.10 in [18], the above covering gives $T(1)$ the structure of a complex Hilbert manifold, modeled on the Hilbert space $A_{2}(\mathbb{D}^{*})$. However, $T(1)$ is not connected with respect to the topology induced by the Hilbert manifold structure.
@@
is a biholomorphic mapping of Hilbert manifolds. As a result, the distribution $\mathcal{D}_{T}$ on $T(1)$ is integrable. The integral manifolds correspond to the components $(\phi+A_{2}(\mathbb{D}^{*}))\cap\beta(T(1))$. For every $[\mu]\in T(1)$, we denote by $T_{[\mu]}(1)$ the component of the Hilbert manifold $T(1)$ containing $[\mu]$. The Hilbert manifold $T_{[\mu]}(1)$ is the integral manifold of the distribution $\mathcal{D}_{T}$ passing through $[\mu]\in T(1)$. In particular, the component of the origin $0\in T(1)$ is denoted by $T_{0}(1)$, and is called the Weil-Petersson universal Teichmüller space.
The Weil-Petersson metric on the Hilbert manifold $T_{0}(1)$ is a Hermitian metric defined by the Hilbert space inner product on tangent space, which is identified with the Hilbert space $H^{-1,1}(\mathbb{D})$ via right translation. Thus, the Weil-Petersson metric is a right invariant metric on $T_{0}(1)$, defined at the origin of $T_{0}(1)$ by
@@
For every $\mu\in H^{-1,1}(\mathbb{D})$, there corresponds a vector field $\frac{\partial}{\partial\epsilon_{\mu}}$ over $V_{0}$. For every $\kappa\in V_{0}$, we define the inner product
@@ 3. Maximal discs of Weil-Petersson class
In this section, we introduce and study maximal discs of Weil-Petersson class in the three-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. We interpret the space of the conformal embeddings of these maximal discs as the cotangent bundle $T^{*}T_{0}(1)$ of the Weil-Petersson universal Teichmüller space $T_{0}(1)$, and show that it can be identified with $T_{0}(1)\times T_{0}(1)$ via the map originally defined by Mess in [12]. For basic terminology related to the geometry of $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we refer to the appendix of this paper. For a more extensive introduction to Anti-de Sitter geometry, we refer to [7].
@@ 3.1. Maximal discs and Gauss maps
We fix the conformal structure on the standard unit disc $\mathbb{D}\subset\mathbb{C}$ and denote it by $\mathbb{D}_{w}$ where the coordinate map $w$ is the identity on $\mathbb{D}$. Given a Beltrami differential $\mu$ representing an element of the universal Teichmüller space $T(1)$, we denote by $\mathbb{D}_{z}$ the corresponding unit disc equipped with the coordinate map $z$, which is the quasi-conformal map $w_{\mu}:\mathbb{D}\to\mathbb{D}$ introduced in the subsection 2.1. Hence, for instance, $w_{\mu}:\mathbb{D}_{z}\to\mathbb{D}_{w}$ is a conformal map, whereas $w_{\mu}:\mathbb{D}_{w}\to\mathbb{D}_{w}$ is not.
For a given $\mathbb{D}_{z}$, we denote its spacelike conformal embedding into the three-dimensional Anti de Sitter space $\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ by
(3.1) $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}.$
The conformal embedding $\sigma$ of the unit disc $\mathbb{D}_{z}$ is called a maximal if the mean curvature identically vanishes, that is, $H_{\sigma}=0$ on the spacelike surface $\Sigma:=\sigma(\mathbb{D})$. The image of a conformal embedding with this property is also refer to as a maximal disc. From now on, we may regard the pullback of a function (or tensor) on $\Sigma$ via $\sigma$ as a function (or tensor) on $\mathbb{D}_{z}$. In particular, we can interpret the mean curvature over $\Sigma$ as a function on $\mathbb{D}_{z}.$
For a spacelike conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, the associated Gauss map, introduced in (A.8), is given by
$G=(G_{+},G_{-}):\mathbb{D}_{z}\to\mathbb{D}_{w}\times\mathbb{D}_{w}.$
By Proposition 3.1 and Theorem 3.3 of [2], we have:
Proposition 3.1.
For a maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, the following hold:
  1. (1)
    Each component of the Gauss map $G_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{w}$ is a harmonic map,
  2. (2)
    The pullback metric of $g_{\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}}$ under $\sigma$, which is the fundamental form $I$, is given by
    $\sigma^{*}(g_{\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}})=I=e^{\phi}|dz|^{2}:=e^{\psi\circ G_{\pm}}\big{|}(G_{\pm})_{z}\big{|}^{2}|dz|^{2},$
    where $e^{\psi(w)}|dw|^{2}$ is the hyperbolic metric on $\mathbb{D}_{w}$.
  3. (3)
    The Hopf differential of $G_{\pm}$ is given by
    $\Phi(G_{\pm})=e^{\psi\circ G_{\pm}}(G_{\pm})_{z}(\overline{G_{\pm}})_{z}\,dz^{2},$
    and it satisfies the relation $\Phi(G_{+})=-\Phi(G_{-})$.
Note that the harmonic map equation for $G_{\pm}$ is given by
(3.2) $(G_{\pm})_{z\bar{z}}+(\psi_{w}\circ G_{\pm})(G_{\pm})_{z}(G_{\pm})_{\bar{z}}=0.$
As evident from this equation, the harmonicity condition on a Riemann surface depends on the metric structure of the target Riemann surface and the conformal structure of the source Riemann surface, rather than the metic structure of the source itself. The underlying fact for Proposition 3.1 is that the pullback metric $G_{\pm}^{*}(e^{\psi(w)}|dw|^{2})$ decomposes as:
(3.3) $G_{\pm}^{*}(e^{\psi(w)}|dw|^{2})=(1+|\mu_{G_{\pm}}|^{2})I+\Phi(G_{\pm})+\overline{\Phi(G_{\pm})},$
where $\mu_{G_{\pm}}$ denotes the Beltrami differential of $G_{\pm}$. The trace part corresponds to $I$ up to the conformal factor $(1+|\mu_{G_{\pm}}|^{2})$, and the off-trace parts are given by the Hopf differential $\Phi(G_{\pm})$, as expected.
The second fundamental form of $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ is given by:
@@
For further details on these equalities, we refer to Section 3 of [2] and Section 5 of [9].
By the Gauss equation for a maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, we obtain the following relation:
@@
Note that this equation can also be derived from the harmonic map equation for $G_{\pm}$ given in (3.2). We also have the following proposition:
@@
The Gaussian curvature $K_{\phi}$ of the metric $I=e^{\phi}|dz|^{2}$ on $\mathbb{D}_{z}$ is given by
@@
For simplicity, let $G=G_{\pm}$ during the proof. By the definition of $e^{\phi}$ and (3.2), we first obtain
@@
$\phi_{z\bar{z}}=(\psi_{w\bar{w}}\circ G)\big{(}|G_{z}|^{2}-|G_{\bar{z}}|^{2}\big{)}.$
Using the Liouville equation for $e^{\psi}$, that is, $\psi_{w\bar{w}}=1/2e^{\psi}$, we get
$\phi_{z\bar{z}}=\frac{1}{2}e^{\psi\circ G}\big{(}|G_{z}|^{2}-|G_{\bar{z}}|^{2}\big{)}.$
Thus, the Gaussian curvature follows as
@@
By the Gauss equation for a maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$, the Gaussian curvature is given by
$K_{\phi}=-1+\kappa^{2}$
where ±$\pm\kappa$ are the principal curvatures of the maximal disc $\Sigma\subset\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$. Consequently, we obtain
@@
The pullback metrics in (3.3) by $G_{\pm}$ are two different hyperbolic metrics, which induce new conformal structures on $\mathbb{D}_{z}$, denoted by $\mathbb{D}_{z_{\pm}}$, respectively. The identity map between $\mathbb{D}_{z}$ and $\mathbb{D}_{z_{\pm}}$ equipped with the hyperbolic metric can then be interpreted as a harmonic map, with its Hopf differential given by ±$\pm\Phi=\Phi(G_{\pm})$, respectively. We denote these maps by
$F_{\pm}:\mathbb{D}_{z}\to\mathbb{D}_{z_{\pm}}.$
Furthermore, by the construction, we have the following:
@@
The metric density $e^{\phi(z)}$ on $\mathbb{D}_{z}$ satisfies the relation
(3.8) $e^{\phi}=e^{\psi\circ G_{\pm}}|(G_{\pm})_{z}|^{2}=e^{\psi\circ F_{\pm}}|(F_{\pm})_{z}|^{2}$
where $e^{\psi}$ represents the hyperbolic metric density on $\mathbb{D}_{w}$, $\mathbb{D}_{z_{\pm}}$, respectively. Moreover, the Beltrami differentials remain invariant, that is,
(3.9) $\mu_{G_{\pm}}=\pm\overline{\Phi}e^{-\phi}=\mu_{F_{\pm}}.$
By Proposition 3.3, the pair $(F_{+},F_{-})$ shares fundamental properties with the Gauss map $(G_{+},G_{-})$. Thus, we refer to $(F_{+},F_{-})$ the induced Gauss map.
Combining the above constructions for both cases of $F_{\pm}$, we obtain the following commutative diagram
@@
where $z,z_{+},z_{-}$ denote the quasi-conformal maps from $\mathbb{D}_{w}$ to $\mathbb{D}_{z}$, $\mathbb{D}_{z_{+}}$, $\mathbb{D}_{z_{-}}$ respectively. Furthermore, we assume that all maps $z,z_{\pm}$ and $F_{\pm}$ are normalized, meaning they preserve the points $1,-1,-\sqrt{-1}$.
By Theorem 4.1 of [2], Theorem 1.10 of [5], and the previous construction, we obtain the following proposition:
@@
For a fixed conformal structure on $\mathbb{D}_{z}$, there exist one-to-one correspondences between the following:
@@
    A maximal conformal embedding $\sigma:\mathbb{D}_{z}\to\mathbb{A}\mathrm{d}\mathbb{S}^{2,1}$ with a complete induced metric and Gaussian curvature that is negative and bounded away from zero.
@@
    An orientation-preserving minimal Lagrangian diffeomorphism $G:\mathbb{D}_{w}\to\mathbb{D}_{w}$ where $\mathbb{D}_{w}$ is equipped with the hyperbolic metric.
@@
    An orientation-preserving minimal Lagrangian diffeomorphism $F:\mathbb{D}_{z_{-}}\to\mathbb{D}_{z_{+}}$ where $\mathbb{D}_{z_{\pm}}$ are equipped with the hyperbolic metric.